# LZ Penalty: An information-theoretic repetition penalty for autoregressive language models

### v1

@@ Abstract
We introduce the LZ penalty, a penalty specialized for reducing degenerate repetitions in autoregressive language models without loss of capability. The penalty is based on the codelengths in the LZ77 universal lossless compression algorithm. Through the lens of the prediction-compression duality, decoding the LZ penalty has the interpretation of sampling from the residual distribution after removing the information that is highly compressible. We demonstrate the LZ penalty enables state-of-the-art open-source reasoning models to operate with greedy (temperature zero) decoding without loss of capability and without instances of degenerate repetition. Both the industry-standard frequency penalty and repetition penalty are ineffective, incurring degenerate repetition rates of up to $4\%$.
@@ 1 Introduction
There are two industry-standard penalties aimed at reducing repetition. First, the repetition penalty Keskar et al. (2019) applies a fixed logit penalty that encourages the model to use new tokens. The frequency penalty, is more subtle, and applies a logit penalty proportional to the token count in context. Neither penalty consistently stops degenerate repetitions without degrading the sample quality. First, the repetition penalty does not actually succeed in preventing degenerate repetitions because it applies a naive, binary modal penalty which does not take into account the number of times a token has appeared. Furthermore, if the repetition penalty is set too high in an effort to minimize this mode collapse, the sampler becomes unable to use fundamental, necessary, but frequent tokens, such as spaces or periods, resulting in poor completions. Thus, merely increasing the repetition penalty is not a viable solution either.
@@
The fundamental improvement in the LZ penalty relative to the repetition or frequency penalty is that the LZ penalty, borrowing from the sliding window matching techniques pioneered in the LZ77 Lempel and Ziv (1977) and LZSS Storer and Szymanski (1982) lossless compression algorithms, depends on the repetition of $n$-grams over a long but fixed-length sliding window. By penalizing as a function of length-normalized $n$-gram statistics as opposed to single token statistics, the penalty can be significantly more surgical in how it modulates the sampling distribution.
@@
We propose the LZ penalty, a novel sampling penalty that demonstrates excellent mitigation of degenerate repetition with greedy sampling in frontier open-source models without loss of capability on difficult reasoning benchmarks such as AIME and GPQA. For practitioners, we provide a batteries-included reference implementation [^2] https://github.com/tginart/sglang in SGLang [^3] https://docs.sglang.ai/. While not fully optimized, the reference implementation incurs a slowdown of less than $0.03\%$ in median throughput and no measurable slowdown in median latency.
@@ 2.1 Language Modeling and Sampling
Language model traces its origins back to Shannon’s testament Shannon (1948), where he trained a casual language model by computing the $n$-gram frequencies over an English text corpus. Modern language models are predominantly based on the transformer architecture Vaswani et al. (2017). Completions from transformer language models are generated by autoregressive sampling of the next-token distribution. The development of transformer language models has been accompanied by significant advancements in sampling techniques that govern text generation. Early explorations of language modeling employed top-$k$ sampling Jozefowicz et al. (2016) to constrain the output distribution to the $k$ most probable tokens, a technique later refined in Welleck et al. (2019), which paired it with unlikelihood training to mitigate repetition. Concurrently, temperature sampling Bowman et al. (2016) emerged as a method to control randomness. Later, nucleus sampling was introduced as a proposed improvement over top-$k$ sampling Holtzman et al. (2020).
These sampling strategies evolved alongside efforts to address text degeneration and repetition. GPT-2’s implementation Radford et al. (2019) implicitly utilized frequency penalties to enhance output fluency, and, concurrently the repetition penalty Keskar et al. (2019) was devised to prevent repetitions and encourage diversity in completions. Later, LaMDA Thoppilan et al. (2022) applied repetition penalties to improve dialog coherence, reflecting a growing emphasis on balancing creativity and quality in LLM outputs. Together, these contributions and others eventually led to an industry-standard sampler which supports a temperature, a top-$k$, a top-$p$ and a frequency or repetition penalty, all of which can be used in tandem to transform the raw next token logits into a final distribution for sampling.
@@
We pause here to define some basic mathematical notation for language modeling.
@@ 2.2 Data Compression
Data compression algorithms go back to the turn of the 20th century. In Shannon’s testament Shannon (1948), he describes the first provably optimal compressor (albeit with exponential time and space). Later, many entropy-optimal compressors achieved practical computational complexity assuming known data distributions. Later still, in 1977 and 1978, the first universal compressors were launched, LZ77 and LZ78, that could, asymptotically, achieve the entropy-rate of any stationary, ergodic data source Lempel and Ziv (1977, 1978); Wyner and Ziv (1994); Morita and Kobayashi (1993). Since then, a whole family of LZ-style compressors has emerged Fiala and Greene (1989); Miller and Wegman (1985); Pavlov (2007); Oberhumer (1997); Yoshizaki (1988); Storer and Szymanski (1982); Welch (1984).
@@
Single-Token Data Compressor. A single-token data compressor, $\mathcal{C}:V\rightarrow\{0,1\}^{*}$ maps literal single-tokens to binary codes. We will assume single-token data compressors are complete prefix codes Cover and Thomas (2006).
@@
In this work, we will only focus on the LZ77 family, which we refer to herein as the LZ sliding window compression algorithm, which uses string matching from a buffer over a lookback window. This contrasts the LZ78 family, which favors tree-style dictionaries. Sliding windows are more convenient for GPUs (for example, by using PyTorch’s Paszke et al. (2019) unfold operation) whereas tree-based dictionary methods are more inherently sequential.
@@
Concretely, LZSS Storer and Szymanski (1982) modifies LZ77 by using a 1-bit flag to indicate whether the next chunk of data is a literal or a length-distance pair and uses literals if a length–distance pair is below a given minimum length. Since we do not actually need to encode or decode the token sequence, the details of the encoding subroutine are not particularly important for our purposes. Instead, we should focus on how many bits are required for the encodings — the codelengths of the resulting codes. We take LZSS as our reference compressor herein, and use the LZSS encoding scheme in the LZ penalty. When we refer to a generic LZ sliding window algorithm, we will mean the LZSS variant.
@@
We define findLongestMatch as the following objective over input strings $y$ and $z$.
$d,l=\arg\max_{j}\left(\max_{k\leq|y|}\left\{k\mid y^{k}=z_{j}^{j+k}\right\}\right)$
@@
$\mathcal{C}_{LZ}(\mathbf{b}|\mathbf{w})=\begin{cases}\mathcal{C}_{LZ}(L,D)&\text{if }L\geq 1\textbf{ and }L=|\mathbf{b|}\\ \mathcal{C}_{LZ}(L,D)\mathcal{C}_{LZ}(\mathbf{b}_{>L}|\mathbf{w})&\text{if }L>1\textbf{ and }L<|\mathbf{b|}\\ \mathcal{C}_{LZ}(\mathbf{b}_{1})\mathcal{C}_{LZ}(\mathbf{b}_{>1}|\mathbf{w})&\text{if }L\leq 0\\ \end{cases}$
@@
Storer and Szymanski (1982) LZSS can encode a match of length $L$ occurring $D$ tokens in the past using $|\mathcal{C}_{LZ}(L,D)|=\log L+\log D+1$ bits.
@@
Storer and Szymanski (1982) LZSS requires $|\mathcal{C}_{LZ}(a)|=\log V+1$ bits to encode token literals $a\in V$.
@@ 2.3 The Prediction-Compression Duality
Then, the Arithmetic coding construction (Cover and Thomas (2006); Witten et al. (1987)) produces a sequential prefix-free compressor $\mathcal{C}_{\mathbf{LM}}$ satisfying, for every $x\in V^{n}$,
@@
Given a language model, we also have a data compressor that compresses as well as the language model predicts, and given a data compressor, we have a language model that predicts as well as that data compressor can compress. Note that the Arithmetic code (and other codes such as the Huffman code Huffman (1952); Cover and Thomas (2006)), employ some version of the prediction compression duality to assign codelengths based on log-probabilities.
However, the situation can be a bit more complex for constructing casual language models form online data compressors such as LZ sliding window algorithms. This is because casual language models must be able to generate a valid next-token pmf at every step whereas data compressors often buffer tokens together into a single code. Practically, this means data compressors don’t necessarily produce a codelength for every next-token. We address this issue by simulating a full buffer and lookback window anew at each next-token rather than exactly implementing the buffer control logic in LZSS. Similar ideas have been explored in Ryabko (2007), which proposes techniques for estimating non-parametric probability densities of online compression algorithms.
@@ 4 Results
Figure 3: Line charts showing the accuracy and repetition percentage for a baseline (repetition penalty of 1, frequency penalty of 0), the LZ penalty, the repetition penalty, and the frequency penalty. Accuracy error bars indicate the empirical std. dev. over 5 runs. We feature the best performing choice of repetition penalty (1.1) and frequency penalty (0.1).
@@ 4.1 Repetition and Capability Benchmarks
We apply the LZ penalty to QwQ-32B [^4] Qwen/QwQ-32B and R1-Distill-14B555deepseek-ai/DeepSeek-R1-Distill-Qwen-14B. We run GPQA and AIME benchmarks (averaging scores and computing std. dev. over 5 runs). We set a max token limit of $24k$. We fixed the top-p to $0.95$ and the top-k to $40$ for all runs. [^6] These are the recommended sampling parameters by the Qwen team For all runs also using the LZ penalty, we fix the penalty strength $\alpha$ to $0.15$, the window size to $512$ and the buffer size to $32$. We found that this configuration of hyperparameters seemed to work well across both models and both datasets with minimal tuning required. [^7] We only tested other penalty strengths in an original sweep of 0.1, 0.2, and 0.3. We found that 0.2 was sometimes too high and 0.1 was sometimes too low. Thus 0.15 seemed to be a sweet spot. The window and buffer sizes were selected on intuition and did not seem to require changing. We detect degenerate repetitions via dual verification of a GPT-4o based judge and a naive search for exact repetitions. [^8] Any substring repeated at least 20 times.
@@
We compare the LZ penalty against two industry-standard penalties: the repetition penalty and the frequency penalty. In both cases, we finely sweep small values up until getting to large values. For the repetition penalty, we sweep $\{1,1.1,1.2,1.25,1.3,1.5\}$ with $1$ being the default value for a no-penalty baseline and $1.2$ being a standard value suggest by Keskar et al. (2019). For the frequency penalty, we sweep $\{0,0.1,0.2,0.3,0.4,0.6,0.8,0.9,1.0\}$ with 00 being the default value for a no-penalty baseline.
@@ Discussion
Based on Fig. 2 and Fig. 3, we observe that neither penalty is a reliable solution. The frequency penalty fails dramatically even for low values such as $0.1$. We suspect that this is because of the length of the generations. Reasoning models produce reasoning traces that can be several thousand tokens long, which simply overwhelms the frequency penalty on common but essential tokens. This often results in a degradation that eventually collapses into a degenerate cycle behavior. The repetition penalty works significantly better than the frequency penalty and does seem to provide some modest relief at low temperatures. However, it is far from a complete solution, with low temperature degenerate repetition rates up to about $\sim 4\%$ depending on model and task domain. This would be disqualifying for any kind of serious application. On the other hand, the LZ penalty achieves effectively zero degenerate repetitions without affecting top-line benchmark scores. The LZ penalty works because it adaptively penalizes based on both the length of the match as well as how far back the match occurs. LZ penalty’s strength increases quickly in match length and attenuates gradually with distance. Both the repetition penalty and the frequency penalty operate at the single-token level whereas the LZ penalty operates at the $n$-gram level. Furthermore, neither the repetition penalty nor the frequency penalty can forget tokens, whereas the LZ penalty quickly and then gradually weakens as the token becomes less recent, until it moves beyond the lookback window altogether and is ignored.
@@ 4.2 Latency and Throughput Benchmark
While our SGLang reference implementation is most certainly not fully optimized, we take minimum care to ensure that it is vectorized (and batched, assuming that the buffer size and lookback size are constant across the batch). [^9] They should be. While we encourage exploration of these hyperparameters, we found buffer size of $32$ and lookback size of $512$ work well with minimal tuning.
We run SGLang’s built-in benchmark script on an 8×8$\times$8 ×H100 node and compare the effect of adding non-zero LZ penalty. While the LZ penalty adds an ultimately negligible amount of computation, it still is significantly more than, say, the repetition penalty, so this it is worthwhile to confirm that we can maintain inference performance.

### v2

@@ Abstract
We introduce the Lemepel-Ziv (LZ) penalty, a penalty specialized for reducing degenerate repetitions in autoregressive language models without loss of capability. The penalty is based on the codelengths in the LZ77 universal lossless compression algorithm. Through the lens of the prediction-compression duality, decoding with the LZ penalty has the interpretation of sampling from the residual distribution after removing the information that is highly compressible. We demonstrate the LZ penalty enables open-source reasoning models to operate with greedy decoding without loss of capability and without instances of degenerate repetition. Both the industry-standard frequency penalty and repetition penalty are ineffective, incurring degenerate repetition rates of up to $4\%$ or more.
@@ 1 Introduction
There are two industry-standard penalties aimed at reducing repetition. First, the repetition penalty [7] applies a fixed logit penalty that encourages the model to use new tokens. The frequency penalty, is more subtle, and applies a logit penalty proportional to the token count in context. Neither penalty consistently stops degenerate repetitions without degrading the sample quality. First, the repetition penalty does not actually succeed in preventing degenerate repetitions because it applies a naive, binary modal penalty which does not take into account the number of times a token has appeared. Furthermore, if the repetition penalty is set too high in an effort to minimize this mode collapse, the sampler becomes unable to use fundamental, necessary, but frequent tokens, such as spaces or periods, resulting in poor completions. Thus, merely increasing the repetition penalty is not a viable solution either.
@@
The fundamental improvement in the LZ penalty relative to the repetition or frequency penalty is that the LZ penalty, borrowing from the sliding window matching techniques pioneered in the LZ77 [8] and LZSS [18] lossless compression algorithms, depends on the repetition of $n$-grams over a long but fixed-length sliding window. By penalizing as a function of length-normalized $n$-gram statistics as opposed to single token statistics, the penalty can be significantly more surgical in how it modulates the sampling distribution.
@@
We propose the LZ penalty, a novel sampling penalty that demonstrates excellent mitigation of degenerate repetition with greedy sampling in frontier open-source models without loss of capability on difficult reasoning benchmarks such as AIME and GPQA. For practitioners, we provide a batteries-included reference implementation [^2] https://github.com/tginart/sglang, https://github.com/tginart/lz-penalty in SGLang. While not fully optimized, the reference implementation incurs a slowdown of less than $0.03\%$ in median throughput and no measurable slowdown in median latency.
Limitations
There are two main limitations to the LZ penalty. First, it is somewhat more complex than the simpler frequency penalty. Second, the LZ penalty addresses determinism at the algorithmic-level, but algorithmic determinism is not always necessary or sufficient in practice. At the application-level, persistent caching or fixed seeding could be used to ensure the same input context consistently produces the same completion. Furthermore, hardware-level non-determinism is possible on certain accelerators, potentially defeating determinism at the algorithmic-level.
@@ 2.1 Language Modeling and Sampling
Language model traces its origins back to Shannon’s testament [17], where he trained a casual language model by computing the $n$-gram frequencies over an English text corpus. Modern language models are predominantly based on the transformer architecture [20]. Completions from transformer language models are generated by autoregressive sampling of the next-token distribution. The development of transformer language models has been accompanied by significant advancements in sampling techniques that govern text generation. Early explorations of language modeling employed top-$k$ sampling [6] to constrain the output distribution to the $k$ most probable tokens, a technique later refined in [22], which paired it with unlikelihood training to mitigate repetition. Concurrently, temperature sampling [1] emerged as a method to control randomness. Later, nucleus sampling was introduced as a proposed improvement over top-$k$ sampling [4].
These sampling strategies evolved alongside efforts to address text degeneration and repetition. GPT-2’s implementation [15] implicitly utilized frequency penalties to enhance output fluency, and, concurrently the repetition penalty [7] was devised to prevent repetitions and encourage diversity in completions. Later, LaMDA [19] applied repetition penalties to improve dialog coherence, reflecting a growing emphasis on balancing creativity and quality in LLM outputs. Together, these contributions and others eventually led to an industry-standard sampler which supports a temperature, a top-$k$, a top-$p$ and a frequency or repetition penalty, all of which can be used in tandem to transform the raw next token logits into a final distribution for sampling.
@@ 2.2 Data Compression
Data compression algorithms go back to the turn of the 20th century. In Shannon’s testament [17], he describes the first provably optimal compressor. Later, many entropy-optimal compressors achieved practical computational complexity assuming known data distributions. Later still, in 1977 and 1978, the first universal compressors were launched, LZ77 and LZ78, that could, asymptotically, achieve the entropy-rate of any stationary, ergodic data source [8, 9, 24, 11]. Since then, a whole family of LZ-style compressors has emerged [3, 10, 14, 12, 25, 18, 21].
@@
Single-Token Data Compressor. A single-token data compressor, $\mathcal{C}:V\rightarrow\{0,1\}^{*}$ maps literal single-tokens to binary codes. We will assume single-token data compressors are complete prefix codes [2].
@@
In this work, we will only focus on the LZ77 family, which we refer to herein as the LZ sliding window compression algorithm, which uses string matching from a buffer over a lookback window. This contrasts the LZ78 family, which favors tree-style dictionaries. Sliding windows are more convenient for GPUs (for example, by using PyTorch’s [13] unfold operation) whereas tree-based dictionary methods are more inherently sequential.
@@
Concretely, LZSS [18] modifies LZ77 by using a 1-bit flag to indicate whether the next chunk of data is a literal or a length-distance pair and uses literals if a length–distance pair is below a given minimum length. Since we do not actually need to encode or decode the token sequence, the details of the encoding subroutine are not particularly important for our purposes. Instead, we should focus on how many bits are required for the encodings — the codelengths of the resulting codes. We take LZSS as our reference compressor herein, and use the LZSS encoding scheme in the LZ penalty. When we refer to a generic LZ sliding window algorithm, we will mean the LZSS variant.
@@
We define findLongestMatch as the following objective over input strings $y$ and $z$. $d,l=\arg\max_{j}\left(\max_{k\leq|y|}\left\{k\mid y_{\leq k}=z_{j}^{j+k}\right\}\right)$
@@
$\mathcal{C}_{LZ}(\mathbf{b}|\mathbf{w})=\begin{cases}\mathcal{C}(d,l)&\text{if}l\geq 1\textbf{ and }l=|\mathbf{b|}\\ \mathcal{C}(d,l)\mathcal{C}_{LZ}(\mathbf{b}_{>l}|\mathbf{w})&\text{if }l\geq 1\textbf{ and }l<|\mathbf{b|}\\ \mathcal{C}(\mathbf{b}_{1})\mathcal{C}_{LZ}(\mathbf{b}_{>1}|\mathbf{w})&\text{if }l=0\\ \end{cases}$
@@
[18] LZSS can encode a match of length $L$ occurring $D$ tokens in the past using $|\mathcal{C}_{LZ}(L,D)|=\log L+\log D+1$ bits.
@@
[18] LZSS requires $|\mathcal{C}_{LZ}(a)|=\log V+1$ bits to encode token literals $a\in V$.
@@ 2.3 The Prediction-Compression Duality
Then, the Arithmetic coding construction ([2, 23]) produces a sequential prefix-free compressor $\mathcal{C}_{\mathbf{LM}}$ satisfying, for every $x\in V^{n}$,
@@
Given a language model, we also have a data compressor that compresses as well as the language model predicts, and given a data compressor, we have a language model that predicts as well as that data compressor can compress. The Arithmetic code (and other codes such as the Huffman code [5, 2]), employ the prediction-compression duality to assign codelengths based on log-probabilities.
The situation is more complex for constructing casual language models from online data compressors such as LZ sliding window algorithms. This is because casual language models must be able to generate a valid next-token pmf at every step whereas data compressors often buffer tokens together into a single code. Practically, this means data compressors do not necessarily produce a codelength for every next-token. We address this issue by simulating a full buffer and lookback window at each next-token. Similar ideas have been explored in [16].
@@ 3 LZ Penalty
Figure 1: An architecture diagram detailing the flow of an autoregressive sampling loop using the LZ penalty.
@@ 4 Results
Figure 2: Line charts showing the accuracy and repetition percentage for a baseline (repetition penalty of 1, frequency penalty of 0), the LZ penalty, the repetition penalty, and the frequency penalty. Accuracy error bars indicate the empirical std. dev. over 5 runs. We feature the best performing choice of repetition penalty and frequency penalty strengths.
@@ 4.1 Repetition and Capability Benchmarks
We apply the LZ penalty to QwQ-32B [^3] Qwen/QwQ-32B and R1-Distill-14B444deepseek-ai/DeepSeek-R1-Distill-Qwen-14B. We run GPQA and AIME benchmarks (averaging scores and computing std. dev. over 5 runs). We set a max token limit of $24k$. We fixed the top-p to $0.95$ and the top-k to $40$ for all runs. [^5] These are the recommended sampling parameters by the Qwen team For all runs also using the LZ penalty, we fix the penalty strength $\alpha$ to $0.15$, the window size to $512$ and the buffer size to $32$. We found that this configuration of hyperparameters seemed to work well across both models and both datasets with minimal tuning required. [^6] We only tested other penalty strengths in an original sweep of 0.1, 0.2, and 0.3. We found that 0.2 was sometimes too high and 0.1 was sometimes too low. Thus 0.15 seemed to be a sweet spot. The window and buffer sizes were selected on intuition and did not seem to require changing. We detect degenerate repetitions via dual verification of a GPT-4o based judge and a naive search for exact repetitions. [^7] Any substring repeated at least 20 times.
@@
We compare the LZ penalty against two industry-standard penalties: the repetition penalty and the frequency penalty. In both cases, we finely sweep small values up until getting to large values. For the results of the full sweep of penalty values, refer to the Appendix.
@@ Discussion
Based on Fig. 2, we observe that neither penalty is a reliable solution. The frequency penalty fails dramatically even for low values. We suspect that this is because of the length of the generations. Reasoning models produce reasoning traces that can be several thousand tokens long, which simply overwhelms the frequency penalty on common but essential tokens. The repetition penalty works significantly better than the frequency penalty and does seem to provide some modest relief. However, it is far from a complete solution, with low temperature degenerate repetition rates up to about $\sim 4\%$ depending on model and task domain. This would be disqualifying for any kind of serious application. On the other hand, the LZ penalty achieves effectively zero degenerate repetitions without affecting top-line benchmark scores. The LZ penalty works because it adaptively penalizes based on both the length of the match as well as how far back the match occurs. LZ penalty’s strength increases quickly in match length and attenuates gradually with distance. Neither the repetition penalty nor the frequency penalty can forget tokens, whereas the LZ penalty quickly and then gradually weakens as the token becomes less recent, until it moves beyond the lookback window altogether.
@@ 4.2 Latency and Throughput Benchmark
While our SGLang reference implementation is not fully optimized, it it is vectorized and batched. We run SGLang’s built-in benchmark script on an 8×8$\times$8 ×H100 node and compare the effect of adding non-zero LZ penalty. While the LZ penalty adds an ultimately negligible amount of computation, it still is significantly more than, say, the repetition penalty, so this it is worthwhile to confirm that we can maintain inference performance.
Appendix A Appendix
A.1 Supplementary Results
As mentioned in the main text, we include sweeps for the penalty strengths for both the repetition and the frequency penalty. For the repetition penalty, we sweep $\{1,1.1,1.2,1.25,1.3,1.5\}$ with $1$ being the default value for a no-penalty baseline and $1.2$ being a standard value suggest by [7]. For the frequency penalty, we sweep $\{0,0.1,0.2,0.3,0.4,0.6,0.8,0.9,1.0\}$ with 00 being the default value for a no-penalty baseline.